The volume of a frustum of a right circular cone is 52π cm*. Its altitude is
3cm and the measure of its lower radius is three times the measure of its
upper radius. the radius of upper base is 1.1547 cm and the slant height of the frustum is 3.7859 cm.
1. Find the lateral area of the frustum.
1
Expert's answer
2020-11-24T11:14:54-0500

Lateral area = (s/3) * (Sum of circumferences)

Where s= slant height = 3.7859

Lower radius = r

Upper radius = R = 1.1547cm

Therefore, r = 3R = 3(1.1547) = 3.4641cm

Circumference of lower radius = 2 * π * r = 2 * π * 3.4641

=21.77cm

Circumference of upper radius = 2 * π * R = 2 * π * 1.1547

= 7.258cm

Therefore, Lateral area = (3.7859/3) * (21.77 + 7.258) = 36.63cm²